Ivan has 400 pounds of mixed nuts that contain no peanuts. How many pounds of peanuts should he put into the mixed nuts so that 20% of the mixture is peanuts?
Answer by richwmiller(17219)

x/(400+x)=1/5
5x=400+x
4x=400
x=100
add 100 pounds of peanuts
or
1*y+0*400=0.2(400+y)
1*y+0=80+0.2y
1*y-0.2y=80-0
0.8y=80
y=80/0.8
y=100